OpenStudy (anonymous):

4/5 6/7 What is x? ---- = ---- 3 1/2 x

OpenStudy (anonymous):

I don't understand the question

OpenStudy (anonymous):

So it is 4/5 / 3 1/2 = 6/7 / x?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Do you have choices?

OpenStudy (anonymous):

I set it up as 4/5x = 7/2 * 6/7 and then you simplify

OpenStudy (anonymous):

Ok, thanks.

OpenStudy (anonymous):

Don't forget to multiply by 5/4 on both sides to get x by itself

OpenStudy (anonymous):

Ok. Thank you so much.